Please help me with these three problems.
1. Find the solution to the equation log(x)8=1/3.
2. Find the approximate solution to the equation 3(1.5^x)+10=280.
3. Find the solution to the equation 2 log(a)3 + log(a)(x-4)= log(a)(x+8).
() below the log
Thank you very much.
Chris

Answer by fractalier(6550)

1) log(x)8 = 1/3
Change this into an exponential form first and then solve, so we get
x^(1/3) = 8
x = 8^3 = 512
2) 3(1.5^x) + 10 = 280
Here isolate the exponential first and get
3(1.5^x) = 270
1.5^x = 90
Now take the log
x*log(1.5) = log(90)
x = log(90) / log(1.5)
3) 2*log(a)3 + log(a)(x-4) = log(a)(x+8)
Combine the left hand side first
log(a) [3^2*(x-4)] = log(a)(x+8) so that
9(x - 4) = x + 8
9x - 36 = x + 8
8x = 44
x = 11/2